OpenStudy (anonymous):

Input in standard form the equation of the given line.The line that passes through (-2, 4) and is parallel to x - 2 y = 6 Input in standard form the equation of the given line.The line that passes through (-2, 4) and is parallel to x - 2 y = 6 @Mathematics

OpenStudy (anonymous):

y-b=m(x-a) You can get the gradient from the equation the line is parallel to as parallel lines means m1=m2 I got x + 2y = 6

OpenStudy (anonymous):

Yeah that is what I got. Lol

OpenStudy (anonymous):

It is not right

OpenStudy (anonymous):

It looks right to me, are you sure it says parallel? If it say perpendicular it would be different.

OpenStudy (anonymous):

yeah I am sure.